Where does the 0.2x come from when this problem is figured to the next step, just can't see it? 2.5x + 2.3(50 - x) = 118.6
Answer by stanbon(48568)

Where does the 0.2x come from when this problem is figured to the next step, just can't see it? 2.5x + 2.3(50 - x) = 118.6
-----------------
Multiply thru by 10 to get rid of the decimals:
25x + 23(50-x) = 1186
25x + 1150 - 23x = 1186
2x = 36
x = 18
===========
Cheers,
Stan H.